OpenStudy (anonymous):

Convert the numeral to a numeral in the base indicated. 16,629 to base 13

OpenStudy (anonymous):

First calculate the powers of 13 until you get something bigger than 16629 13^0 = 1 13^1 = 13 13^2 = 169 13^3 = 2197 13^4 = 28561....bigger than 16629..so stop. we won't need 13^4

OpenStudy (anonymous):

Now look at 13^3 which is 2197 16629 / 2197 = 7 remainder 1250 1250 / 169 = 7 remainder 67 67 / 13 = 5 remainder 2 2 / 1 = 2 Answer 7752

OpenStudy (unklerhaukus):

if you want to check it \[7752_{13}\\\quad=7\times13^3+7\times13^2+5\times13^1+2\times13^0\\\quad=7\times2197+7\times169+5\times13+2\\\quad=15379+1183+65+2\\\quad=\color{white}{16629}\]